# Dragging with Geometry: From Pixels to Geometry-Guided Image Editing

### v1

@@ 4.3 Ablation Study
Displacement Field. We conduct ablation studies to investigate the contribution of each component in our hybrid displacement field. Specifically, w/o Depth removes the depth-aware field, while w/o Plane removes the plane-aware field. As shown in Fig. 8, removing the depth-aware field leads to inaccurate editing (e.g., failure to rotate the car). Removing plane-aware field leads to insufficient editing. The results highlight the complementary roles of 3D geometry and 2D plane prior in achieving structure-preserving and semantically coherent editing.
Conflict-Free Partitioning. We evaluate the proposed conflict-free partitioning strategy with direct summation (Directly Add), pixel-distance weighting (Pixel Distance), and drag-magnitude weighting (Drag Magnitude). As shown in Fig. 6, directly adding displacement fields leads to conflict and cancellation when directions oppose; Pixel Distance and Drag Magnitude cannot effectively separate influence regions, producing duplicated and unsatisfactory results; Our method avoids interference among multiple drag points, yielding accurate results.
Hyperparameters. We visually analyze the influence of . As shown in Fig. 9(left), setting in the range of 0.5 to 1.5 achieves a good balance between geometry-aware consistency and local editability. We analyze the impact of and in Fig. 9(right). The lower indicates a smoother scale (see Eq. (7)), weakening the influence of depth contrast. This results in larger overall deformation and improved alignment. Larger increases depth sensitivity, causing sharper displacement changes. This allows for finer-grained control in regions with significant depth changes. For plane-aware field, a higher enforces more localized and sharper deformations (see Eq. (8)), thereby better aligning the edits with user intention and resulting in lower MD. Additional ablation results on the noise-scaling term in Eq. 2 are provided in Appendix C.5.
@@ 5 Conclusion
Limitations. Although GeoDrag supports one-step editing, computing multiple displacement fields introduces additional computational overhead compared to lightweight 2D-only baselines such as FastDrag [18]. But GeoDrag often yields much more geometry-consistent and high-quality image edits, e.g., and improvements in terms of MD and GPU memory than FastDrag as shown in Table 1.
@@ Appendix B Implementation Details
GeoDrag is implemented based on a pretrained latent diffusion model (Stable Diffusion 1.5 [26]), and incorporates an LCM-accelerated U-Net [20] to enable efficient low-step inference. We set the number of sampling steps to 10 and use an inversion strength of . The depth prediction model used in geometry-aware field modeling (Sec. 3.1) is Depth Anything V2 [33]. Following prior drag-based editing methods [3, 13, 12], we disable classifier-free guidance. For fair comparison, all baseline methods are evaluated using their default hyperparameter settings as specified in the original papers or official open-source implementations. Experiments are conducted on an RTX 4090 GPU with 24G memory.
@@ Appendix C Supplementary Ablation Study
C.1 Quantitative Ablations on Displacement Field and Multi-Point Dragging
We present additional quantitative results to validate the effectiveness of our hybrid displacement field and conflict-free partitioning strategy. As shown in Fig. 10(a), GeoDrag with geometry- and plane-aware displacement field achieves the best performance. Removing the geometry prior (w/o Depth) or the plane modulation (w/o Plane) leads to degradation across all metrics.
The quantitative results of different multi-point dragging strategies are shown in Fig. 10(b). GeoDrag outperforms alternative strategies on all metrics, demonstrating the effectiveness of the conflict-free partitioning. The mathematical formulations of alternative strategies are provided below:
@@ C.4 Ablation study on inversion steps
In contrast, when , the editing quality degrades. For a fair comparison with FastDrag [18], and to balance editing performance with fidelity to the original image, we adopt 10 as the default setting for inversion steps throughout all experiments in the paper.
C.5 Ablation on the Noise-Scaling Term
We conduct an ablation study on the noise-scaling term in Eq. (2), which is controlled by the stochasticity parameter :
Recall that sets the magnitude of the random noise injected by the masked stochastic DDIM step (Eq. 2). As shown in Fig. 14, setting leads to a fully deterministic update, which often produces over-smoothed results after interpolation and thus blurs in the background. Increasing injects stronger randomness inside the interpolated region, encouraging the diffusion model to better infer local details there. These results highlight the importance of controlled noise injection: moderate stochasticity yields sharper and more faithful refinements without incurring additional sampling overhead.
@@ Appendix E Effect of LoRA Finetuning on GeoDrag
To further investigate the generalization ability of GeoDrag, we evaluate GeoDrag under two configurations: with and without LoRA [34] finetuning. Both versions share the same backbone and inference hyperparameters; the only difference is whether LoRA finetuning is applied. As shown in visual examples (see Fig. 19), GeoDrag consistently produces high-quality, geometry-consistent edits even without any finetuning. Nevertheless, LoRA finetuning can enhance local detail and fidelity in some cases (e.g., the fourth row of Fig. 19).
@@ Appendix G Effect of Masks
Mask selection is crucial for controlling the scope and locality of edits. Figure 16 illustrates the impact of different masks when edit a same object. Applying a full-object mask to the avocado enlarges the entire object (Figure 16(a)) and the shape of the seed is changed. In contrast, masking only the avocado’s edges confines the effect to the boundary (Figure 16(b)), preserving the shape of seed.
@@ Appendix I Details of User Study
Here, we provide additional details about the user study, including its design and aggregated ranking-based evaluation results. Fig. 20 shows the selected input images, the corresponding editing results from different methods, and the user-assigned rankings for each result. For each test case, users were asked to rank the edited results from three different methods. The ranking is based on how well each result aligns with the intended dragging operation while preserving the original visual identity. A lower rank indicates better quality. Specifically, rank 1 denotes the best edit—i.e., the one that best aligns with the drag intention and maintains high image fidelity—while rank 3 corresponds to the least satisfactory result. Note that the anonymous labels (a), (b), and (c) correspond to FastDrag [18], FreeDrag [13], and GeoDrag, respectively.

### v2

@@ 4.3 Ablation Study
Displacement Field. We conduct ablation studies to investigate the contribution of each component in our hybrid displacement field. Specifically, w/o Depth removes the depth-aware field, while w/o Plane removes the plane-aware field. As shown in Fig. 8, removing the depth-aware field leads to inaccurate editing (e.g., failure to rotate the car). Removing plane-aware field leads to insufficient editing. The results highlight the complementary roles of 3D geometry and 2D plane prior in achieving structure-preserving and semantically coherent editing. The quantitative resuls are shown in Fig. 10(a). GeoDrag with geometry- and plane-aware displacement field achieves the best performance. Removing the geometry prior (w/o Depth) or the plane modulation (w/o Plane) leads to degradation across all metrics.
Conflict-Free Partitioning. We evaluate the proposed conflict-free partitioning strategy with direct summation (Directly Add), pixel-distance weighting (Pixel Distance), and drag-magnitude weighting (Drag Magnitude). The mathematical formulations of alternative strategies are provided in Appendix C.1. As shown in Fig. 6, directly adding displacement fields leads to conflict and cancellation when directions oppose; Pixel Distance and Drag Magnitude cannot effectively separate influence regions, producing duplicated and unsatisfactory results; Our method avoids interference among multiple drag points, yielding accurate results. The quantitative results are shown in Fig. 10(b). GeoDrag outperforms alternative strategies on all metrics, demonstrating the effectiveness of the conflict-free partitioning.
Hyperparameters. We visually analyze the influence of . As shown in Fig. 9(left), setting between 0.5 and 1.5 strikes a good balance between geometry-aware consistency and local editability. We analyze the impact of and in Fig. 9(right). The lower indicates a smoother scale (see Eq. (7)), weakening the influence of depth contrast. This results in larger overall deformation and improved alignment. Larger increases depth sensitivity, causing sharper displacement changes. This allows for finer-grained control in regions with significant depth changes. For plane-aware field, a higher enforces more localized and sharper deformations (see Eq. (8)), thereby better aligning the edits with user intention and resulting in lower MD. The study of the noise-scaling term is shown in Fig. 11. Setting leads to a fully deterministic update, often causing over-smoothed interpolations and blurred backgrounds. Increasing adds randomness to the interpolated region, helping the diffusion model better recover local details. More ablation studies and detailed analyses are provided in Appendix C.
@@ 5 Conclusion
Limitations. Although GeoDrag supports one-step editing, computing multiple displacement fields introduces additional computational overhead compared to lightweight 2D-only baselines such as FastDrag (Zhao et al., 2024). But GeoDrag often yields much more geometry-consistent and high-quality image edits, e.g., and improvements in terms of MD and GPU memory than FastDrag as shown in Table 1.
Acknowledgments
This research is supported by the National Natural Science Foundation of China (Nos.52441503 and 62302093), the Natural Science Foundation of Jiangsu Province (Nos.BK20230833), the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (Proposal ID: 23-SIS-SMU-070), and the Big Data Computing Center of Southeast University. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
Ethics statement
Image editing model may contain biases or occasionally produce sensitive or offensive outputs. Our models are presented strictly for academic and scientific research purposes. Any generated content does not reflect the personal views of the authors. Our work remains guided by a commitment to advancing AI technologies in ways that uphold ethical standards and resonate with societal values.
Reproducibility statement
Appendix
@@ Appendix B Implementation Details
GeoDrag is implemented based on a pretrained latent diffusion model (Stable Diffusion 1.5 (Rombach et al., 2022)), and incorporates an LCM-accelerated U-Net (Luo et al., 2023) to enable efficient low-step inference. We set the number of sampling steps to 10 and use an inversion strength of . The depth prediction model used in geometry-aware field modeling (Sec. 3.1) is Depth Anything V2 (Yang et al., 2024a). Following prior drag-based editing methods (Shi et al., 2024; Ling et al., 2024; Liu et al., 2024), we disable classifier-free guidance. For fair comparison, all baseline methods are evaluated using their default hyperparameter settings as specified in the original papers or official open-source implementations. Experiments are conducted on an RTX 4090 GPU with 24G memory.
@@ Appendix C Supplementary Ablation Study
C.1 Mathematical Formulations of Alternative Strategies
@@ C.4 Ablation study on inversion steps
In contrast, when , the editing quality degrades. For a fair comparison with FastDrag (Zhao et al., 2024), and to balance editing performance with fidelity to the original image, we adopt 10 as the default setting for inversion steps throughout all experiments in the paper.
@@ Appendix E Effect of LoRA Finetuning on GeoDrag
To further investigate the generalization ability of GeoDrag, we evaluate GeoDrag under two configurations: with and without LoRA (Hu et al., 2022) finetuning. Both versions share the same backbone and inference hyperparameters; the only difference is whether LoRA finetuning is applied. As shown in visual examples (see Fig. 24), GeoDrag consistently produces high-quality, geometry-consistent edits even without any finetuning. Nevertheless, LoRA finetuning can enhance local detail and fidelity in some cases (e.g., the fourth row of Fig. 24).
@@ Appendix G Effect of Masks
Mask selection is crucial for controlling the scope and locality of edits. Fig. 16 illustrates the impact of different masks when edit a same object. Applying a full-object mask to the avocado enlarges the entire object (Fig. 16(a)) and the shape of the seed is changed. In contrast, masking only the avocado’s edges confines the effect to the boundary (Fig. 16(b)), preserving the shape of seed.
@@ Appendix I Details of User Study
Here, we provide additional details about the user study, including its design and aggregated ranking-based evaluation results. Fig. 25 shows the selected input images, the corresponding editing results from different methods, and the user-assigned rankings for each result. For each test case, users were asked to rank the edited results from three different methods. The ranking is based on how well each result aligns with the intended dragging operation while preserving the original visual identity. A lower rank indicates better quality. Specifically, rank 1 denotes the best edit—i.e., the one that best aligns with the drag intention and maintains high image fidelity—while rank 3 corresponds to the least satisfactory result. Note that the anonymous labels (a), (b), and (c) correspond to FastDrag (Zhao et al., 2024), FreeDrag (Ling et al., 2024), and GeoDrag, respectively.
Appendix J Robustness to Depth Map
| Noisy Level | MD | DAI1 | DAI10 | DAI20 | IF |
| Baseline | 29.24 | 0.128 | 0.120 | 0.111 | 0.847 |
| +0.66 | +0.00 | +0.00 | +0.00 | +0.00 | |
| +1.23 | +0.002 | +0.002 | +0.001 | +0.00 | |
| +2.85 | +0.004 | +0.003 | +0.003 | -0.002 | |
| +9.81 | +0.007 | +0.004 | +0.004 | -0.006 |
Fig. 17 presents a qualitative evaluation of our method under challenging depth-prediction conditions, including transport-specular regions (first row), textureless regions (second row), and a strong perspective view (third row). For each scenario, we compare two variants: geometry-only and the combined formulation employed in GeoDrag. As shown in the figure, the underlying depth maps in these scenes exhibit distinct failure modes: (1) due to the transparent glass surface and strong specular reflections, the depth estimator can only separate foreground and background coarsely, producing a flat and overly simplified depth map. When relying solely on depth cues, this leads to blurred deformation along object boundaries; (2) the edited object lies almost flush against a uniformly colored wall, resulting in an extremely weak depth gradient. Geometry-only guidance, therefore, becomes unreliable and introduces distortions; (3) the strong perspective foreshortening causes large depth discontinuities and makes it difficult for the estimator to recover the true geometry of the target object (the train).
In all these scenarios, the failure of the depth map to provide reliable geometric cues renders geometry-only guidance insufficient for stable deformation. However, the incorporation of Spatial Plane Modulation effectively compensates for these deficiencies by supplying a depth-independent, spatially smooth prior that preserves coherent structure even when depth information is severely degraded. This design enables GeoDrag to maintain stable and plausible deformations.
Moreover, we conducted an additional perturbation study to evaluate the robustness of GeoDrag under noisy monocular depth predictions. We injected controlled Gaussian noise into the predicted depth maps, scaling the variance according to the depth range of each image:
where is the original depth, denotes the depth range, and is the noise factor. This formulation ensures that the injected perturbation is proportional to the scene’s depth variation. As reported in Table 4, GeoDrag exhibits strong robustness to depth prediction noise. Under mild and moderate perturbations (e.g., ), the performance remains largely stable across all evaluation metrics, indicating that the model is not overly sensitive to small depth fluctuations. Even when the noise magnitude increases to , the performance degradation only becomes noticeable at this extreme level, suggesting that the method is capable of maintaining reliable tracking quality under realistic sensor and prediction uncertainties. Overall, these results demonstrate that GeoDrag preserves its effectiveness in challenging noisy-depth scenarios and does not rely heavily on precise depth estimates to function correctly.
Appendix K Failure Cases
We present two representative failure cases in Fig. 18 to clarify the limitations of GeoDrag. Both examples require generating previously unseen content rather than simply deforming the visible regions. The example on the left additionally involves a complex scene with multiple overlapping objects. Although the editing direction is partially correct, the newly exposed regions exhibit distorted and blurry structures. This occurs because our interpolation strategy reconstructs unseen content by sampling nearby latent features; when multiple objects heavily overlap, the mixed semantic context introduces feature ambiguity and leads to unstable reconstruction. The example on the right exposes a large area that is completely invisible in the original image. In this case, the missing content has no reliable correspondence within the existing latent featureslacks reliable correspondence within the existing latent features, so interpolation-based completion cannot provide meaningful structural cues, resulting inpainting-based latent completion module may alleviate this issue and is a promising direction for future work.
Appendix L Per-Stage Runtime Analysis and Improvements
| Stage | DDIM inversion | DDIM sampling | Interpolation | Depth prediction |
| Time (s) | 1.2721 | 1.2416 | 0.2247 | 0.0995 |
| Stage | Spatial plane modulation | Latent relocation | Partition | Geometry-aware field modeling |
| Time (s) | 0.0109 | 0.2714 | 0.0006 | 0.0003 |
| Model | MD | DAI1 | DAI10 | DAI20 | IF | Interpolation Time (s) | Latent Relocation Time (s) |
| GeoDrag (original) | 29.24 | 0.128 | 0.120 | 0.111 | 0.847 | 0.2247 | 0.2714 |
| GeoDrag (optimized) | 27.84 | 0.133 | 0.121 | 0.111 | 0.847 | 0.0014 | 0.0009 |
To better characterize the efficiency of the proposed GeoDrag pipeline, we provide a full-stack runtime breakdown across all modules, revealing where the computational budget is primarily spent and how to further optimized. The pipeline can be decomposed into eight major stages: DDIM inversion, DDIM sampling, partition, geometry-aware field modeling, spatial plane modulation, interpolation, latent relocation, and depth prediction. The averaged per-stage execution times (in seconds) over the Other Object subset of DragBench are reported in Table 5. The results show that the diffusion-based components (inversion and sampling) dominate the total latency, whereas the remaining geometric modules contribute negligibly to the runtime. Within the non-diffusion stages, interpolation and latent relocation are most expensive component.
Therefore, we focuse our efficiency improvements specifically on these two modules: (1) we observed that latent relocation can be fully vectorized, allowing us to remove unnecessary per-point loops; and (2) we replace the dynamic BNNI update with a static four-neighbor interpolation rule, enabling parallel computation and substantially faster execution. The quantitative comparison is reported in Table 6. Surprisingly, this simplified interpolation not only reduces runtime but also improves performance. We believe this is because removing the iterative BNNI updates prevents error accumulation, while the static strategy still retains the essential semantic structure in the latent space.
Appendix M Effect of the Scaling Factor on Handle–Target Alignment
During the drag operation, the displacement field is computed in the diffusion latent space rather than directly on image pixels space. When converting drag points from the pixel domain (e.g., pixels) to the latent domain (e.g., features), the coordinates must be downscaled accordingly. This conversion inevitably reduces the effective displacement magnitude in the latent space, leading to a small residual gap between the handle and target points after editing.
To compensate for this attenuation, we apply a scaling factor to rescale the displacement field after converting coordinates to the latent space. This restores the motion magnitude lost during downsampling and improves handle–target alignment.
As shown in Fig. 19, the scaling factor significantly influences alignment and deformation quality. A moderate value enhances correspondence between handle and target points, while an overly large one leads to overcorrection and geometric distortion. In practice, a scale of offers a good balance between point-level accuracy and global consistency.
Appendix N Additional Visual Comparisons with SDEDrag and GoodDrag
We provide additional qualitative comparisons with SDEDrag (Nie et al., 2024) and GoodDrag (Zhang et al., 2025) in Fig. 20. The results show that our method achieves comparable or even better editing results. Compared with GoodDrag (Zhang et al., 2025), GeoDrag does not rely on LoRA fine-tuning and multi-step optimization, making it significantly more efficient.
Declaration of LLM usage
We used a large language model (GPT-5) solely for grammar and language refinement. All research ideas, analyses, and conclusions are our own.